# A Contextual Quality Reward Model for Reliable and Efficient Best-of-N Sampling

### v1

@@ Abstract
Keywords Reward Model $\cdot$ Inference Time Alignment $\cdot$ Discrete Choice
@@ 1 Introduction
The capabilities of Large Language Models (LLMs) have been significantly enhanced through alignment with human feedback—making them more helpful and harmless (Bai et al., 2022), instruction-following (Ouyang et al., 2022), and capable of sophisticated tasks (Ziegler et al., 2019). Such advances are driven by preference-based alignment methods, ranging from inference-time techniques like Best-of-N (BoN) sampling (Nakano et al., 2021; Ouyang et al., 2022) to post-training optimization approaches like Direct Preference Optimization (DPO) (Rafailov et al., 2023). Crucially, the effectiveness of these approaches depends entirely on the data used to train their underlying preference models.
Current methods predominantly rely on pairwise comparison datasets, where humans select the better of two responses. This paradigm, rooted in models like Bradley–Terry (Bradley and Terry, 1952), has an inherent limitation: it only captures relative preference, not response acceptability. While the field often implicitly treats the chosen response as “desirable” (Jung et al., 2024; Ethayarajh et al., 2024), the data merely indicates that it is better than the alternative. As a result, a reward model trained this way can rank responses but cannot determine whether a response is acceptable or unacceptable (Wang et al., 2023). This limitation critically undermines BoN sampling. There is no guarantee that the best of $N$ options meets a minimum quality threshold. It is possible that BoN simply selects the least bad of many poor options.
@@
First, we provide a rigorous analysis and empirical evidence showing that standard Best-of-N (BoN) sampling is unreliable for challenging prompts. We demonstrate that as more candidates are generated, the risk of a false acceptance-selecting a response that is merely the least bad and does not meet a minimum quality threshold—increases significantly. Our experiments show that the number of such reliability failures for the baseline method more than doubles when scaling from $N=1$ to $N=32$.
@@
For reliability critical applications, the framework can be configured as what we term an Alignment Guardrail. By setting a calibrated quality threshold, the system refrains from outputting a response unless it is demonstrably acceptable. This is essential in contexts like customer-facing chatbots. For instance, rather than providing a policy-compliant but unhelpful response (the least bad option it could generate), the system can recognize that no candidate will satisfy the user and instead escalate the query to a human agent. This prevents user frustration and ensures a higher standard of interaction. Our experiments confirm this is highly effective, reducing the number of false acceptances by 70% compared to the standard BoN baseline. This ability to abstain or invoke fallback strategies is a crucial capability for safe and reliable systems (Lightman et al., 2023; Snell et al., 2024; Kang et al., 2025; Bai et al., 2022).
@@
This establishes a clear and tunable trade-off between reliability and computational efficiency that is absent in current methods.
The remainder of this paper is structured as follows. Section 2 reviews related work in preference alignment and inference efficiency. Section 3 details the choice-theoretic formulation of our reward model. Section 4 presents the framework for estimating the model’s parameters and evaluating its performance. Section 5 introduces our primary methodological contribution, the best of mini-N in-loop strategy, and develops the mechanics for its two configurations: the alignment guardrail and the inference accelerator. Section 6 provides experimental results, first confirming the failure modes of standard BoN and then demonstrating the practical effectiveness of our two configurations. Finally, Section 7 concludes and discusses future work.
@@ 2 Related Work
Most modern LLM alignment techniques, from RLHF to DPO, are trained on datasets of pairwise preferences (Ouyang et al., 2022; Rafailov et al., 2023). While effective for learning relative ranking, this paradigm fails to capture a signal of contextual acceptability. The concept of using an outside option to enrich this data has been explored by Wang et al. (2024). However, in that work, the outside option was used as a data cleaning heuristic; any sample where the annotator chose to reject all responses was simply discarded from the dataset. Similarly, Rho (2025) investigated using an outside option for post-training alignment methods like DPO and RLHF. Their analysis suggests that while the outside option is not directly applicable to the DPO setup and does not alter RLHF outcomes, removing samples where the outside option was chosen can improve DPO performance—further treating the signal as a data cleaning tool.
Our work departs from prior approaches by treating the outside option not as a data cleaning heuristic, but as a core learning signal. We use this signal within a discrete choice framework to explicitly teach the reward model the concept of a contextual “acceptability threshold.” This allows the model to distinguish not just what is better, but what is good enough—a critical capability for reliable inference-time control.
BoN sampling is a powerful but computationally expensive method for improving response quality at inference time. Several recent works have focused on mitigating its high cost. For example, speculative rejection sampling aims to halt the generation of unpromising candidates midway through (Sun et al., 2024), while Adaptive BoN seeks to dynamically select a smaller $N$ for easier prompts (Raman et al., 2025). Our best of mini-N in-loop method offers a complementary strategy. Rather than pruning unpromising candidates, our approach leverages its signal of contextual acceptability to actively identify the first certifiably acceptable candidate and terminate the process early. This provides a statistically grounded stopping criterion that other efficiency-oriented methods lack. While we do not explore this here, our early-exit strategy could likely be composed with methods like speculative rejection sampling or Adaptive BoN for further efficiency gains, representing a promising direction for future work.
@@ 3 A Choice-Based Reward Model
In this section, we lay out the theoretical formulation of our choice-based reward model. We begin by adapting the McFadden discrete choice model (McFadden, 2001), a standard framework from economics, to include an “outside option” representing a labeller’s choice to reject all candidate responses. We then specify the utility model and derive the multinomial logit choice probabilities. From this, we define a normalized reward function that is anchored to the outside option. We then analyze the properties of this function, showing that it preserves relative preference rankings while providing an interpretable signal of contextual acceptability, as the underlying normalized reward can be uniquely identified from choice data.
@@ 3.2.2 Identification of Contextual Acceptability
This can be seen by rearranging Eq. (2) to express the normalized reward directly in terms of choice probabilities:
@@ 4.1 Log-Likelihood Function
By substituting the choice probability from Eq. (2), the log-likelihood for a single observation where the labeller chose option $d_{k}$ is given by:
@@
Note that the ‘1’ in the denominator of the log term corresponds to $\exp(R_{\theta}(x_{k},y_{k,0}))$, as the normalized reward of the outside option is zero by definition. The model is trained by maximizing the total log-likelihood in Eq. (3) using a standard gradient-based optimizer.
@@ 4.2.1 Data Transformation for Binary Evaluation
It is important to note that this data transformation uses the observed human choice as a proxy for the true, latent quality of a response. Ideally, the ground truth label for a response $y$ would be determined by the sign of its true reward, i.e., $R(x,y)>0$, as this represents the average preference across a population of labellers, not just the stochastic utility of a single individual. However, due to the stochastic nature of the choice model specification (1), we only observe a noisy realization of this preference. The metrics are therefore conditioned on the observed choice, $d_{k}$, rather than the unobserved reward sign. For a given threshold $\tau$ (which is 0 in our framework), we are essentially using the labeller’s choice to approximate the ideal but unobservable target. Instead of directly measuring a quantity like $\Pr(\hat{R}_{\theta}(x,y)>\tau|R(x,y)>0)$, our metrics are based on the observable proxy $\Pr(\hat{R}_{\theta}(x,y)>\tau|\text{response }y\text{ was chosen})$. This is the best available approximation of the desired target.
@@ 4.2.2 Probabilistic Model and Metrics
Given this binary dataset and the choice model specification (1), we can assess the model’s performance. The probability of a response $y$ being acceptable is modeled as $P(1|x,y)=\sigma(\hat{R}_{\theta}(x,y))$, where $\sigma(\cdot)$ is the sigmoid function. The probability of it being unacceptable is $P(0|x,y)=1-\sigma(\hat{R}_{\theta}(x,y))=\sigma(-\hat{R}_{\theta}(x,y))$.
@@ 4.3 A Note on Class Imbalance and Choice-Based Sampling
Attempting to correct for this imbalance via re-sampling techniques introduces a form of sampling bias. As demonstrated by Manski and Lerman (1977), estimating a choice model on a dataset where the choice proportions have been artificially altered leads to inconsistent estimates of the underlying reward function. Specifically, if one uses an altered sampling distribution $H(i)$ instead of the true population distribution $Q(i)$ to estimate the model with the maximum likelihood objective in Eq. (3), the resulting reward function, $R^{\prime}$, becomes a biased estimator of the true reward $R$:
@@ 5.1 The Failure Mode of Standard BoN
We analyze the terms in this expression to determine the sign of the difference. The first term, $\sum q_{N}(k)[a_{k,N+1}-a_{k,N}]$, is also non-negative due to the monotonicity of $a_{k,N}$ as established in Eq. (4). The sign of the final term depends on the difference $q_{N}(k)-q_{N}(k-1)$, which is positive when $k<(N+1)p_{g}$ and negative when $k>(N+1)p_{g}$.
@@ 6.1.2 Training Procedure
    Finally, we simulate a human labeller’s choice by adding i.i.d. Gumbel noise to the ground-truth scores of $y_{1}$, $y_{2}$, and the outside option (with a score of 0). The option with the highest noisy score is recorded as the choice, $d\in\{0,1,2\}$ as illustrated by Eq. (1). This process yields a dataset of tuples: $(x,y_{1},y_{2},d)$.
@@
    Reward Model Training: The final reward model is trained on 10,000 samples from this synthetic preference dataset, with 2,000 samples held out for evaluation. The model is trained to maximize the log-likelihood as defined in Eq. (3).
@@ 7 Conclusion
Appendix A Acknowledgements
We used OpenAI ChatGPT, Google Gemini, and xAI Grok for literature search support, wording/clarity edits, mathematical cross-checks, and code review/debugging. The authors verified all outputs and are solely responsible for the paper’s content.
The experiments presented in this paper were made possible by the computational resources provided by Google Colab and the Kaggle platform.
Appendix B Best of Mini N in Loop Algorithm
@@
1:Input: Prompt $x$, total budget NN, mini-batch size $n$, number of loops $L=N/n$, tuning mode $m\in\{\text{'Alignment Guardrail'},\text{'Inference Accelerator'}\}$.
@@
6:if $m=\text{'Alignment Guardrail'}$ then

### v2

# Best of mini-N in-loop Sampling: A Contextual Quality Reward Model for Reliable and Efficient Best-of-N Sampling

@@ 1 Introduction
The capabilities of Large Language Models (LLMs) have been significantly enhanced through alignment with human feedback, making them more helpful and harmless (Bai et al., 2022), instruction-following (Ouyang et al., 2022), and capable of sophisticated tasks (Ziegler et al., 2019). Such advances are driven by preference-based alignment methods, ranging from inference-time techniques like Best-of-N (BoN) sampling (Nakano et al., 2021; Ouyang et al., 2022) to post-training optimization approaches like Direct Preference Optimization (DPO) (Rafailov et al., 2023). Crucially, the effectiveness of these approaches depends entirely on the data used to train their underlying preference models.
Current methods predominantly rely on pairwise comparison datasets, where humans select the better of two responses. This paradigm, rooted in models like Bradley–Terry (Bradley & Terry, 1952), has an inherent limitation: it only captures relative preference, not response acceptability. While the field often implicitly treats the chosen response as “desirable” (Jung et al., 2024; Ethayarajh et al., 2024), the data merely indicates that it is better than the alternative. As a result, a reward model trained this way can rank responses but cannot determine whether a response is acceptable or unacceptable (Wang et al., 2023). This limitation critically undermines BoN sampling. There is no guarantee that the best of $N$ options meets a minimum quality threshold. It is possible that BoN simply selects the least bad of many poor options.
@@
First, we present a theoretical analysis and empirical evidence showing that standard Best-of-N (BoN) sampling is increasingly unreliable for challenging prompts. As $N$ grows, the probability of a false acceptance (i.e., selecting the least-bad candidate that still fails a minimum quality threshold) rises substantially. Our experiment results show that the number of such reliability failures in the BoN baseline more than doubles as $N$ increase from 1 to 32.
@@
For reliability critical applications, the framework can be configured as what we term an Alignment Guardrail. By setting a calibrated quality threshold, the system refrains from outputting a response unless it is demonstrably acceptable. This is essential in contexts like customer-facing chatbots. For instance, rather than providing a policy-compliant but unhelpful response (the least bad option it could generate), the system can recognize that no candidate will satisfy the user and instead escalate the query to a human agent. This prevents user frustration and ensures a higher standard of interaction. Our experiments confirm this is highly effective, reducing the number of false acceptances by 70% compared to the standard BoN baseline. This ability to abstain or invoke fallback strategies is a crucial capability for safe and reliable systems (Lightman et al., 2024; Snell et al., 2024; Kang et al., 2025; Bai et al., 2022).
@@
This establishes a clear and tunable trade-off between reliability and computational efficiency that is absent in prior methods.
@@ 2 Related Work
Modern LLM alignment techniques, most prominently Reinforcement Learning from Human Feedback (RLHF) and RL-free direct methods like Direct Preference Optimization (DPO), are trained on datasets of pairwise preferences (Ouyang et al., 2022; Rafailov et al., 2023). Such data is inherently ordinal, so reward models can rank candidates, but lack an absolute acceptability notion, which allows “least-bad” selections on hard prompts. The notion of incorporating an outside option into preference data has been previously noted by Wang et al. (2024). In particular, the HelpSteer2-Preference dataset introduces a “Neither response is valid” label, functioning as an implicit outside option when both candidate responses are unacceptable. However, this signal was employed solely as a data-cleaning heuristic, samples with this label were excluded during preprocessing rather than integrated into the reward modeling process.
Our work departs from prior approaches by treating the outside option not as a data cleaning heuristic, but as a core learning signal. We use this signal within a discrete choice framework to explicitly teach the reward model the concept of a contextual “acceptability threshold.” This allows the model to distinguish not just what is better, but what is good enough, a critical capability for reliable inference-time control.
BoN is an effective but compute-intensive approach to improving inference-time quality. Inference-time selection methods, including BoN and reward-model reranking, raise output quality without modifying the base policy, as demonstrated by rejection sampling against a reward model in WebGPT (Nakano et al., 2021) and by broader alignment experiments on preference-modeling and scaling (Askell et al., 2021). Nevertheless, when all samples are weak, these methods can still return the least-bad candidate. To reduce computational burden, recent work prunes low-promise generations mid-decode via speculative rejection sampling (Sun et al., 2024), adapts the sample size $N$ to prompt difficulty in Adaptive BoN (Raman et al., 2025), introduces confidence-based early exit for generation (Schuster et al., 2022), and accelerates decoding through speculative decoding while preserving the output distribution (Leviathan et al., 2023).
Our best-of-mini-$N$ in-loop strategy is complementary: instead of pruning partial candidates or fixing $N$ in advance, it sequences small batches and terminates as soon as a candidate exceeds a context-calibrated acceptability threshold, yielding a principled, quality-based stopping rule. Although we leave integration to future work, the procedure is naturally compatible with pruning, adaptive $N$, early-exit, and speculative decoding, and can be combined with these techniques for additional efficiency gains.
@@ 3 A Choice-Based Reward Model
We develop a choice-based reward model by adapting McFadden’s discrete-choice model (McFadden, 2001) to include an explicit “outside option” that captures a labeler’s decision to reject all candidates. We then specify the utility model and derive the resulting multinomial logit choice probabilities. From these probabilities, we construct a normalized reward anchored to the outside option and analyze its properties: the reward preserves relative preference rankings while supplying an interpretable, context-dependent acceptability signal. Under standard identification conditions, this normalized reward is uniquely recoverable from choice data.
@@ 3.2.2 Identification of Contextual Acceptability
This can be seen by rearranging Eq. equation 2 to express the normalized reward directly in terms of choice probabilities:
@@ 4.1 Log-Likelihood Function
By substituting the choice probability from Eq. equation 2, the log-likelihood for a single observation where the labeller chose option $d_{k}$ is given by:
@@
Note that the ‘1’ in the denominator of the log term corresponds to $\exp(R_{\theta}(x_{k},y_{k,0}))$, as the normalized reward of the outside option is zero by definition. The model is trained by maximizing the total log-likelihood in Eq. equation 3 using a standard gradient-based optimizer.
@@ 4.2.1 Data Transformation for Binary Evaluation
It is important to note that this data transformation uses the observed human choice as a proxy for the true, latent quality of a response. Ideally, the ground truth label for a response $y$ would be determined by the sign of its true reward, i.e., $R(x,y)>0$, as this represents the average preference across a population of labellers, not just the stochastic utility of a single individual. However, due to the stochastic nature of the choice model specification equation 1, we only observe a noisy realization of this preference. The metrics are therefore conditioned on the observed choice, $d_{k}$, rather than the unobserved reward sign. For a given threshold $\tau$ (which is 0 in our framework), we are essentially using the labeller’s choice to approximate the ideal but unobservable target. Instead of directly measuring a quantity like $\Pr(\hat{R}_{\theta}(x,y)>\tau|R(x,y)>0)$, our metrics are based on the observable proxy $\Pr(\hat{R}_{\theta}(x,y)>\tau|\text{response }y\text{ was chosen})$. This is the best available approximation of the desired target.
@@ 4.2.2 Probabilistic Model and Metrics
Given this binary dataset and the choice model specification equation 1, we can assess the model’s performance. The probability of a response $y$ being acceptable is modeled as $P(1|x,y)=\sigma(\hat{R}_{\theta}(x,y))$, where $\sigma(\cdot)$ is the sigmoid function. The probability of it being unacceptable is $P(0|x,y)=1-\sigma(\hat{R}_{\theta}(x,y))=\sigma(-\hat{R}_{\theta}(x,y))$.
@@ 4.3 A Note on Class Imbalance and Choice-Based Sampling
Attempting to correct for this imbalance via re-sampling techniques introduces a form of sampling bias. As demonstrated by Manski & Lerman (1977), estimating a choice model on a dataset where the choice proportions have been artificially altered leads to inconsistent estimates of the underlying reward function. Specifically, if one uses an altered sampling distribution $H(i)$ instead of the true population distribution $Q(i)$ to estimate the model with the maximum likelihood objective in Eq. equation 3, the resulting reward function, $R^{\prime}$, becomes a biased estimator of the true reward $R$:
@@ 5.1 The Failure Mode of Standard BoN
We analyze the terms in this expression to determine the sign of the difference. The first term, $\sum q_{N}(k)[a_{k,N+1}-a_{k,N}]$, is also non-negative due to the monotonicity of $a_{k,N}$ as established in Eq. equation 4. The sign of the final term depends on the difference $q_{N}(k)-q_{N}(k-1)$, which is positive when $k<(N+1)p_{g}$ and negative when $k>(N+1)p_{g}$.
@@ 6.1.2 Training Procedure
    Finally, we simulate a human labeller’s choice by adding i.i.d. Gumbel noise to the ground-truth scores of $y_{1}$, $y_{2}$, and the outside option (with a score of 0). The option with the highest noisy score is recorded as the choice, $d\in\{0,1,2\}$ as illustrated by Eq. equation 1. This process yields a dataset of tuples: $(x,y_{1},y_{2},d)$.
@@
    Reward Model Training: The final reward model is trained on 10,000 samples from this synthetic preference dataset, with 2,000 samples held out for evaluation. The model is trained to maximize the log-likelihood as defined in Eq. equation 3.
@@ 7 Conclusion
Appendix A Appendix: Best of Mini N in Loop Algorithm
@@
1:Input: Prompt $x$, total budget NN, mini-batch size $n$, number of loops $L=N/n$, tuning mode $m\in\{\text{`Alignment Guardrail'},\text{`Inference Accelerator'}\}$.
@@
6:if $m=\text{`Alignment Guardrail'}$ then